# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 5.1. Kostant-Mickelsson-Whittaker algebras
$I_{k}:=(A^{(j)}(\Omega_{j})|j=1\ldots k)\cap\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}$
the ideal inside $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}$. For any $k$, there is an isomorphism
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\cong\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}/I_{k}.$
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\cong\bigoplus_{k=0}^{\infty}\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}/I_{k}.$
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\subset\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+}$
@@
$\mathbf{Z}_{\mathfrak{gl}_{n}}[\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}\rightarrow\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]/(A^{(j)}(\Omega_{j})|j=1,\ldots,k)$
is equal to $S_{k}$-invariants of the target. However, it can be identified with $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}/I_{k}$ by definition. ∎
The following theorem relates the positive part of the shifted Yangian from ˜2.13 to the Kostant-Mickelsson-Whittaker algebra associated to $S^{\bullet}\mathbf{C}^{N}$. Observe similarity of the explicit formulas below to the quantization of Zastava spaces for $\mathrm{SL}_{2}$ constructed in [FR14].
@@
$x_{+}(u)$ $\mapsto-\frac{1}{u-\Omega_{1}}(v_{1})=-\sum_{i=0}^{\infty}\Omega_{1}^{i}(v_{1})u^{-i-1}\in(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+})[[u^{-1}]].$
@@
$A(z)\frac{1}{w-\Omega_{1}}(v_{1})+\frac{1}{w-\Omega_{1}}(v_{1})A(z)=-\frac{\hbar}{z-w}A(z)\left(\frac{1}{z-\Omega_{1}}-\frac{1}{w-\Omega_{1}}\right)(v_{1}).$
@@
and, due to Proposition˜4.7, the left-hand side is
$A(z)\frac{1}{w-\Omega_{1}}(v_{1})+A(z)\frac{1}{w-\Omega_{1}}(v_{1})\left(1+\frac{\hbar}{z-\Omega_{1}}\right)=A(z)\frac{\hbar}{(z-\Omega_{1})(w-\Omega_{1})}(v_{1}).$
@@
As for the relation (2.12), observe that due to the second part of Proposition˜5.7 and ˜5.8, the product
$(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+})\otimes(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{l}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+})\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k+l}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$
@@
$\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}\otimes\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{l}]^{S_{l}}$ $\rightarrow\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k+l}]^{S_{k+l}},$
$f(\Omega_{1},\ldots,\Omega_{k})\otimes g(\Omega_{1},\ldots,\Omega_{l})$ $\mapsto e^{\hbar}_{k+l}\cdot f(\Omega_{1},\ldots,\Omega_{k})g(\Omega_{k+1},\ldots,\Omega_{k+l}),$
where $e^{\hbar}_{k+l}$ is the symmetrizer of Section˜2.2. However, upon extending the scalars to $\mathbf{Z}_{\mathfrak{gl}_{N}}$, this is exactly the product of the shuffle algebra $\mathcal{F}^{\hbar}$ of ˜2.8, in other words, the “quotient” map
$\mathcal{F}^{\hbar}=\bigoplus_{k=0}^{\infty}\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}\rightarrow\bigoplus_{k=0}^{\infty}\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}/I_{k}\cong\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$
is an algebra map. By Theorem˜2.18, there is an algebra isomorphism $Y(\mathfrak{gl}_{2})^{+}\xrightarrow{\sim}\mathcal{F}^{\hbar}$. Twisting the resulting composition $Y(\mathfrak{gl}_{2})^{+}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ by negative sign (which does not influence the defining relations) sends $x_{+}(u)$ to $\frac{1}{u-\Omega_{1}}(v_{1})$. ∎
5.2. Other representations
In this subsection, we compare the isomorphism of Proposition˜5.7 with the trivialization (5.3) provided by the Kirillov projector.
@@ 5.2. Other representations
$\varphi\colon\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes\mathbf{C}^{N}\otimes\ldots\otimes\mathbf{C}^{N}\rightarrow\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]/(A^{(j)}(\Omega_{j})|j=1,\ldots,k)$
of left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-modules is completely determined by two conditions: for every multi-index $\{N\geq i_{1}\geq\ldots\geq i_{k}\geq 1\}$, we have
$\varphi(v_{i_{1}}\otimes\ldots\otimes v_{i_{k}})=\psi^{k-i_{1}-\ldots-i_{k}}\Omega_{1}^{i_{1}-1}\ldots\Omega^{i_{k}-1}_{k},$
@@
$\varphi(v_{i_{\sigma(1)}}\otimes\ldots\otimes v_{i_{\sigma(k)}})=\sigma^{\hbar}\cdot\varphi(v_{i_{1}}\otimes\ldots\otimes v_{i_{k}}),$
where the action is that of the degenerate affine Hecke algebra.
@@
In what follows, we will drop the tensor product sign by considering $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$ as a submodule of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\bigoplus_{i=0}^{\infty}(\mathbf{C}^{N})^{\otimes i}$ and using notations of ˜5.6.
@@
$P_{\psi}v_{i_{1}}P_{\psi}\ldots P_{\psi}v_{i_{k}}P_{\psi}=P_{\psi}v_{i_{1}}\ldots v_{i_{k}}P_{\psi}$
for sequences $\{i_{1}\geq\ldots\geq i_{k}\}$ since
$P_{\psi}v_{i_{1}}P_{\psi}\ldots P_{\psi}v_{i_{k}}P_{\psi}=\psi^{k-i_{1}-\ldots-i_{k}}P_{\psi}\Omega_{1}^{i_{1}-1}\ldots\Omega_{k}^{i_{k}-1}(v_{1}^{\otimes k})P_{\psi}.$
We will prove it by induction on each term of the sequence $\{i_{1}\geq\ldots\geq i_{k}\}$. The base is
$P_{\psi}v_{1}P_{\psi}\ldots P_{\psi}v_{1}P_{\psi}=P_{\psi}v_{1}\ldots v_{1}P_{\psi},$
@@
$P_{\psi}v_{i_{1}}P_{\psi}\ldots P_{\psi}v_{i_{k}}P_{\psi}=P_{\psi}v_{i_{1}}\ldots v_{i_{k}}P_{\psi}.$
Take $1\leq\alpha\leq k$ such that $i_{\alpha-1}\geq i_{\alpha}+1\geq i_{\alpha+1}$ (i.e. the non-increasing condition is still satisfied). By ˜5.4, we have
$P_{\psi}v_{i_{1}}P_{\psi}\ldots P_{\psi}v_{i_{\alpha}+1}P_{\psi}\ldots P_{\psi}v_{i_{k}}P_{\psi}=\psi^{-1}P_{\psi}v_{i_{1}}P_{\psi}\ldots P_{\psi}\Omega_{\alpha}(v_{i_{\alpha}})P_{\psi}\ldots P_{\psi}v_{i_{k}}P_{\psi}.$
@@
$P_{\psi}v_{i_{1}}\ldots\Omega_{\alpha}(v_{i_{\alpha}})\ldots v_{i_{k}}P_{\psi}=P_{\psi}v_{i_{1}}\ldots v_{i_{\alpha-1}}(\sum_{\beta=1}^{N}E_{i_{\alpha}\beta}v_{\beta})v_{i_{\alpha+1}}\ldots v_{i_{k}}P_{\psi}.$
Consider the sum for $\beta>i_{\alpha}$. Observe that due to $\{i_{1}\leq\ldots\leq i_{\alpha}\}$, the terms $E_{i_{\alpha}\beta}$ commute with the head $v_{i_{\alpha+1}}\ldots v_{i_{k}}$, hence we have
$P_{\psi}v_{i_{1}}\ldots v_{i_{\alpha-1}}(\sum_{\beta>i_{\alpha}}E_{i_{\alpha}\beta}v_{\beta})v_{i_{\alpha+1}}\ldots v_{i_{k}}P_{\psi}=$
$=$ $P_{\psi}v_{i_{1}}\ldots v_{i_{\alpha-1}}(\sum_{\beta>i_{\alpha}}v_{\beta}v_{i_{\alpha+1}}\ldots v_{i_{k}}E_{i_{\alpha}\beta})P_{\psi}+\hbar(N-i_{\alpha})P_{\psi}v_{i_{1}}\ldots v_{i_{\alpha}}\ldots v_{i_{k}}P_{\psi}=$
$=$ $\psi P_{\psi}v_{i_{1}}\ldots v_{i_{\alpha}+1}\ldots v_{i_{k}}P_{\psi}+\hbar(N-i_{\alpha})P_{\psi}v_{i_{1}}\ldots v_{i_{\alpha}}\ldots v_{i_{k}}P_{\psi}$
due to the properties of Theorem˜5.2 of the Kirillov projector. Consider the sum for $\beta\leq i_{\alpha}$. Since we assumed that $i_{1}\geq\ldots\geq i_{\alpha-1}>i_{\alpha}$, the elements $E_{i_{\alpha}\beta}$ also commute with the tail $v_{i_{1}}\ldots v_{i_{\alpha-1}}$. Using again the properties of the Kirillov projector, we have
$P_{\psi}v_{i_{1}}\ldots v_{i_{\alpha-1}}(\sum_{\beta\leq i_{\alpha}}E_{i_{\alpha}\beta}v_{\beta})v_{i_{\alpha+1}}\ldots v_{i_{k}}P_{\psi}=-\hbar(N-i_{\alpha})P_{\psi}v_{i_{1}}\ldots v_{i_{\alpha}}\ldots v_{i_{k}}P_{\psi}.$
@@
$P_{\psi}v_{i_{1}}P_{\psi}\ldots P_{\psi}v_{i_{\alpha}+1}P_{\psi}\ldots P_{\psi}v_{i_{k}}P_{\psi}=P_{\psi}v_{i_{1}}\ldots v_{i_{\alpha}+1}\ldots v_{i_{k}}P_{\psi},$
@@
Therefore, for every irreducible representation $V$ with an embedding $V\hookrightarrow(\mathbf{C}^{N})^{\otimes k}$, there is an explicit realization of the corresponding Kostant-Whittaker reduction
$\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes V\cong\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes V/\!\!/_{\psi}\,N_{+}\rightarrow\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]/(A^{(j)}(\Omega_{j})|j=1\ldots k).$
@@
$V[\hbar]\rightarrow\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]_{<N},$
@@
5.3. Mirabolic reduction and Feigin-Odesskii shuffle algebras

### v4

@@ 6.1 Kostant–Mickelsson–Whittaker algebras
$I_{k}:=\bigl(A^{(j)}(\Omega_{j})\mid j=1,\dots,k\bigr)\cap\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]^{S_{k}}$
the ideal inside $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]^{S_{k}}$. For any $k$, there is an isomorphism
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\cong\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]^{S_{k}}/I_{k}.$
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\cong\bigoplus_{k=0}^{\infty}\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]^{S_{k}}/I_{k}.$
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\subset\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+}$
@@
$\mathbf{Z}_{\mathfrak{gl}_{n}}[\Omega_{1},\dots,\Omega_{k}]^{S_{k}}\rightarrow\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]/\bigl(A^{(j)}(\Omega_{j})\mid j=1,\dots,k\bigr)$
is equal to $S_{k}$-invariants of the target. However, it can be identified with $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]^{S_{k}}/I_{k}$ by definition. ∎
The following theorem relates the positive part of the shifted Yangian from Definition 3.13 to the Kostant–Mickelsson–Whittaker algebra associated to $S^{\bullet}\mathbf{C}^{N}$. Observe similarity of the explicit formulas below to the quantization of Zastava spaces for $\mathrm{SL}_{2}$ constructed in [15].
@@
$x_{+}(u)$ $\mapsto-\frac{1}{u-\Omega_{1}}(v_{1})=-\sum_{i=0}^{\infty}\Omega_{1}^{i}(v_{1})u^{-i-1}\in\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\bigr)\bigl[\hskip-0.56905pt\bigl[u^{-1}\bigr]\hskip-0.56905pt\bigr].$
@@
$A(z)\frac{1}{w-\Omega_{1}}(v_{1})+\frac{1}{w-\Omega_{1}}(v_{1})A(z)=-\frac{\hbar}{z-w}A(z)\biggl(\frac{1}{z-\Omega_{1}}-\frac{1}{w-\Omega_{1}}\biggr)(v_{1}).$
@@
and, due to Proposition 5.7, the left-hand side is
$A(z)\frac{1}{w-\Omega_{1}}(v_{1})+A(z)\frac{1}{w-\Omega_{1}}(v_{1})\biggl(1+\frac{\hbar}{z-\Omega_{1}}\biggr)=A(z)\frac{\hbar}{(z-\Omega_{1})(w-\Omega_{1})}(v_{1}).$
@@
As for the relation (3.12), observe that due to the second part of Proposition 6.7 and Corollary 6.8, the product
$\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\bigr)\otimes\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{l}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\bigr)\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k+l}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$
@@
$\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]^{S_{k}}\otimes\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{l}]^{S_{l}}$ $\rightarrow\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k+l}]^{S_{k+l}},$
$f(\Omega_{1},\dots,\Omega_{k})\otimes g(\Omega_{1},\dots,\Omega_{l})$ $\mapsto{\rm e}^{\hbar}_{k+l}\cdot f(\Omega_{1},\dots,\Omega_{k})g(\Omega_{k+1},\dots,\Omega_{k+l}),$
where ${\rm e}^{\hbar}_{k+l}$ is the symmetrizer of Section 3.2. However, upon extending the scalars to $\mathbf{Z}_{\mathfrak{gl}_{N}}$, this is exactly the product of the shuffle algebra $\mathcal{F}^{\hbar}$ of Definition 3.8, in other words, the “quotient” map
$\mathcal{F}^{\hbar}=\bigoplus_{k=0}^{\infty}\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]^{S_{k}}\rightarrow\bigoplus_{k=0}^{\infty}\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]^{S_{k}}/I_{k}\cong\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$
is an algebra map. By Theorem 3.18, there is an algebra isomorphism $Y(\mathfrak{gl}_{2})^{+}\xrightarrow{\sim}\mathcal{F}^{\hbar}$. Twisting the resulting composition $Y(\mathfrak{gl}_{2})^{+}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ by negative sign (which does not influence the defining relations) sends $x_{+}(u)$ to $\frac{1}{u-\Omega_{1}}(v_{1})$. ∎
6.2 Other representations
In this subsection, we compare the isomorphism of Proposition 6.7 with the trivialization (6.3) provided by the Kirillov projector.
@@ 6.2 Other representations
$\varphi\colon\ \mathbf{Z}_{\mathfrak{gl}_{N}}\otimes\mathbf{C}^{N}\otimes\dots\otimes\mathbf{C}^{N}\rightarrow\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]/\bigl(A^{(j)}(\Omega_{j})\mid j=1,\dots,k\bigr)$
of left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-modules is determined by two conditions: for every multi-index $\{N\geq i_{1}\geq\dots\geq i_{k}\geq 1\}$, we have
$\varphi(v_{i_{1}}\otimes\dots\otimes v_{i_{k}})=\psi^{k-i_{1}-\dots-i_{k}}\Omega_{1}^{i_{1}-1}\cdots\Omega^{i_{k}-1}_{k},$
@@
$\varphi(v_{i_{\sigma(1)}}\otimes\dots\otimes v_{i_{\sigma(k)}})=\sigma^{\hbar}\cdot\varphi(v_{i_{1}}\otimes\dots\otimes v_{i_{k}}),$
where the action on the right is that of the degenerate affine Hecke algebra.
@@
In what follows, we will drop the tensor product sign by considering $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$ as a submodule of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\bigoplus_{i=0}^{\infty}\bigl(\mathbf{C}^{N}\bigr)^{\otimes i}$ and using notations of Definition 6.6.
@@
$P_{\psi}v_{i_{1}}P_{\psi}\cdots P_{\psi}v_{i_{k}}P_{\psi}=P_{\psi}v_{i_{1}}\cdots v_{i_{k}}P_{\psi}$
for sequences $\{i_{1}\geq\dots\geq i_{k}\}$ since
$P_{\psi}v_{i_{1}}P_{\psi}\cdots P_{\psi}v_{i_{k}}P_{\psi}=\psi^{k-i_{1}-\dots-i_{k}}P_{\psi}\Omega_{1}^{i_{1}-1}\cdots\Omega_{k}^{i_{k}-1}\bigl(v_{1}^{\otimes k}\bigr)P_{\psi}.$
We will prove it by induction on each term of the sequence $\{i_{1}\geq\dots\geq i_{k}\}$. The base is
$P_{\psi}v_{1}P_{\psi}\cdots P_{\psi}v_{1}P_{\psi}=P_{\psi}v_{1}\cdots v_{1}P_{\psi},$
@@
$P_{\psi}v_{i_{1}}P_{\psi}\cdots P_{\psi}v_{i_{k}}P_{\psi}=P_{\psi}v_{i_{1}}\cdots v_{i_{k}}P_{\psi}.$
Take $1\leq\alpha\leq k$ such that $i_{\alpha-1}\geq i_{\alpha}+1\geq i_{\alpha+1}$ (i.e., the non-increasing condition is still satisfied). By Lemma 6.4, we have
$P_{\psi}v_{i_{1}}P_{\psi}\cdots P_{\psi}v_{i_{\alpha}+1}P_{\psi}\cdots P_{\psi}v_{i_{k}}P_{\psi}=\psi^{-1}P_{\psi}v_{i_{1}}P_{\psi}\cdots P_{\psi}\Omega_{\alpha}(v_{i_{\alpha}})P_{\psi}\cdots P_{\psi}v_{i_{k}}P_{\psi}.$
@@
$P_{\psi}v_{i_{1}}\cdots\Omega_{\alpha}(v_{i_{\alpha}})\cdots v_{i_{k}}P_{\psi}=P_{\psi}v_{i_{1}}\cdots v_{i_{\alpha-1}}\Biggl(\sum_{\beta=1}^{N}E_{i_{\alpha}\beta}v_{\beta}\Biggr)v_{i_{\alpha+1}}\cdots v_{i_{k}}P_{\psi}.$
Consider the sum for $\beta>i_{\alpha}$. Observe that due to $\{i_{1}\leq\dots\leq i_{\alpha}\}$, the terms $E_{i_{\alpha}\beta}$ commute with the head $v_{i_{\alpha+1}}\cdots v_{i_{k}}$, hence we have
$P_{\psi}v_{i_{1}}\cdots v_{i_{\alpha-1}}\biggl(\sum_{\beta>i_{\alpha}}E_{i_{\alpha}\beta}v_{\beta}\biggr)v_{i_{\alpha+1}}\cdots v_{i_{k}}P_{\psi}$
$\qquad{}=P_{\psi}v_{i_{1}}\cdots v_{i_{\alpha-1}}\biggl(\sum_{\beta>i_{\alpha}}v_{\beta}v_{i_{\alpha+1}}\cdots v_{i_{k}}E_{i_{\alpha}\beta}\biggr)P_{\psi}+\hbar(N-i_{\alpha})P_{\psi}v_{i_{1}}\cdots v_{i_{\alpha}}\cdots v_{i_{k}}P_{\psi}$
$\qquad{}=\psi P_{\psi}v_{i_{1}}\cdots v_{i_{\alpha}+1}\cdots v_{i_{k}}P_{\psi}+\hbar(N-i_{\alpha})P_{\psi}v_{i_{1}}\cdots v_{i_{\alpha}}\cdots v_{i_{k}}P_{\psi}$
due to the properties of (6.2) of the Kirillov projector. Consider the sum for $\beta\leq i_{\alpha}$. Since we assumed that $i_{1}\geq\dots\geq i_{\alpha-1}>i_{\alpha}$, the elements $E_{i_{\alpha}\beta}$ also commute with the tail $v_{i_{1}}\cdots v_{i_{\alpha-1}}$. Using again the properties of the Kirillov projector, we have
$P_{\psi}v_{i_{1}}\cdots v_{i_{\alpha-1}}\biggl(\sum_{\beta\leq i_{\alpha}}E_{i_{\alpha}\beta}v_{\beta}\biggr)v_{i_{\alpha+1}}\cdots v_{i_{k}}P_{\psi}=-\hbar(N-i_{\alpha})P_{\psi}v_{i_{1}}\cdots v_{i_{\alpha}}\cdots v_{i_{k}}P_{\psi}.$
@@
$P_{\psi}v_{i_{1}}P_{\psi}\cdots P_{\psi}v_{i_{\alpha}+1}P_{\psi}\cdots P_{\psi}v_{i_{k}}P_{\psi}=P_{\psi}v_{i_{1}}\cdots v_{i_{\alpha}+1}\cdots v_{i_{k}}P_{\psi},$
@@
Therefore, for every irreducible representation $V$ with an embedding $V\hookrightarrow\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$, there is an explicit realization of the corresponding Kostant–Whittaker reduction
$\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes V\cong\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes V/\!\!/_{\psi}\,N_{+}\rightarrow\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]/\bigl(A^{(j)}(\Omega_{j})\mid j=1,\dots,k\bigr).$
@@
$V[\hbar]\rightarrow\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]_{<N},$
@@
6.3 Mirabolic reduction and Feigin–Odesskii shuffle algebras